### v1

Abstract
The category was introduced in [12] in order to provide a structural setting to the study of the many Gelfand-Tsetlin subalgebras appearing in the context of the diagrammatic Soergel category of Elias and Williamson [3]. The category have as objects, all that we called nil graded algebras associated to a triangular matrix and as morphisms, all the preserving degree homomorphisms of graded algebras between them. In this article we develop a series of isomorphism criteria on that we used later to find out relevant information on the Gelfand-Tsetlin subalgebras of the diagrammatic Soergel category.
Contents
1 Introduction
1.1 Motivation
The diagrammatic Soergel Category was defined in 2016 by Elias and Williamson [3] in order to provide a diagrammatic approach to the study of the category of Soergel bimodules, a well known incarnation of the Hecke category, [18]. Since then, the category have gained attention of both algebraists and algebraic geometrist, due to its intrinsic connection with both branches of study.
A particular interest is the study of the internal structure of the category itself, that is, the structure of the many Hom spaces and more particularly of the many End algebras. In this context, for example, the diagrammatic setting provided by Elias and Williamson have been crucial to find out and prove the cellularity of the Light leaves bases of Libedinsky [8],[3]. Another example of the power of this diagrammatic approach is the discovery due to Ryom-Hansen, of explicit Jucys-Murphy elements relative to the Libedinsky’s bases [19].
In [10], Libedinsky and Plaza, found out a hard connection between the category and what they called the Blob category, another diagrammatic category, inspired in the well known KLR-algebras of Khovanov, Lauda and Rouquier. The similar patterns that they observed in both categories motivate them to define the Blob vs Soergel conjecture, where roughly speaking, they declare that there is an explicit incarnation of the category (in type ), in terms of KLR-diagrams.
Some of the first evidences on the veracity of the aforementioned conjecture can be found in the work of Espinoza and Plaza [4], where explicit isomorphisms were built between the corresponding Gelfand-Tsetlin subalgebras, that is, the algebras generated by the Jucys-Murphy elements, coming on one side from Two-color Soergel Calculus and on the other of certain idempotent truncation of the KLR version of the blob algebra (see [17]). In [13], Plaza, Ryom-Hansen and the author, obtain isomorphisms in a higher level, this time between certain endomorphism algebras of in type and the corresponding idempotent truncation of blob algebras. In both cases the procedure implies the search of optimal presentations in terms of generators and relations for the algebras involved, therefore, in both [4] and [13], the reader will find, not only a clear proof of each isomorphism, but also deep information on the structure of the corresponding algebras.
Recently in [1], Bowman, Cox and Hazi have proved the Blob vs Soergel conjecture, they in fact, proved a more general result, obtaining as a consequence the equivalences predicted by Libedinsky and Plaza. The smart procedure used in [1], based in alcove geometry and paths configurations, allows the authors to cover all the cases in just one go, while in [4] and [13] only particular cases were covered. On the other hand the general approach developed in [1] do not implies the aforementioned internal information on the corresponding algebras obtained in [4] and [13] or any generalization of that. Therefore, we consider as an open problem the search of this internal information for all the cases no covered in [4] and [13], that is in principle, to find out optimal presentations for all the algebras related with the Blob vs Soergel conjecture.
In the last few years, the author have been developing certain strategies to address this problem. For example in [11], presentations for the Gelfand-Tsetlin subalgebras of a family of idempotent truncations of the generalized blob algebras were found, later in [12], working on the side of the category the author developed a series of techniques and procedures to obtain optimal presentation for the Gelfand-Tsetlin subalgebras of the many End algebras for any type (far beyond the initial objective: type ). Both [11] and [12] are natural generalizations of [4] and particularly they seem to imply a hard connection between the light leaves bases of Libedinsky and the graded cellular bases built by Ryom-Hansen and the author in [14], just as was previously observed in [4] and [13], in that case, with respect to the graded cellular bases developed by Plaza and Ryom-Hansen in [17].
The category was introduced in [12] in order to provide a structural setting to the study of the many Gelfand-Tsetlin subalgebras appearing in the context of the diagrammatic Soergel category The category have as objects, all that we called nil graded algebras associated to a triangular matrix and as morphisms, all the preserving degree homomorphisms of graded algebras between them. Naturally, the aforementioned Gelfand-Tsetlin subalgebras of are (the main) objects in our category, and all the knowledge of developed in [12] allowed us to find out the desired presentations for those algebras. On the other hand, the category is by itself an interesting object of study and the results obtained in [12] open the door to a deeper analysis on it.
Being more specific, the objects of the category are commutative graded algebras defined by a presentation with a finite set of generators and certain quadratic relations codified by a strictly triangular matrix. In [12] we define and make a first study of those algebras, obtaining a series of techniques and corresponding applications to the study of the many Gelfand-Tsetlin subalgebras of the category Since each objects of is totally determined by a codifying triangular matrix, it seem plausible to ask when two of those objects are isomorphic and which role play the corresponding codifying matrices in this aspect. To answer this question is the main motivation of this and other works we have in progress.
In section 2, we develop a series of isomorphism criteria on that partially answer the above question. In Theorem 2.2 we obtain an absolute (but not easy to apply) matrix criterion to determine if two objects are or not isomorphic, this is the key result that allowed us to find out different approaches to a complete classification of the objects of Applying conveniently theorem 2.2, we could describe completely the class of the algebra associated to the zero matrix obtaining a very clear and easy to apply criterium (Theorem 2.9) to determine if the algebra is isomorphic to or not, depending on the entries of the codifying matrix The case of the matrix although very particular, can be cleverly used to determine if other two objects of are or not isomorphic (see for example the proof of proposition 3.3). Another fundamental result of this article is Theorem 2.39, where we describe another absolute criterium (again, not always easy to apply) to determine if two objects are or not isomorphic. In the case of theorem 2.39, the criterion states that two objects and are isomorphic if and only if, there is a finite sequence of certain matrix operations (that we called elementary triangular operations), that progressively transform the matrix into the matrix
Our first examples are developed in subsection 2.2, where the smaller cases were studied. The results that we found there inspire all the analysis we did in this article. In particular, they open interesting questions for further works, for example, since we observe that for small matrices, we have a very small number of isomorphic classes in then is natural to ask:
- Q1:
Is the number of classes always finite?
Although we do not have an answer for this question yet, in subsection 2.4, we obtain a lower bound for the number of classes depending on the size of the codifying matrices. In theorem 2.23, we describe an explicit set of representatives of different classes, for each size We left Q1 open, together with the search of a complete set of representatives and explicit criteria to describe any isomorphic class such as we did in the case of the matrix
On the other hand, in section 3, we advocate our attention to apply the results developed in section 2, to the study of the Gelfand-Tsetlin subalgebras coming from the diagrammatic Soergel category In [12], for each element we a define the SLTM, and the corresponding nil graded algebra We have proved in [12], that the algebra corresponds to the Gelfand-Tsetlin subalgebra of the endomorphism algebra in the diagrammatic Soergel Category of Elias and Williamson, relative to the Libedynski’s light leaves basis and the set of Jucys-Murphy elements defined by Ryom-Hansen (See [3],[8] and [19]).
In [12] we show that for the case of a Coxeter system of type the algebras are invariant under commuting moves. That result naturally open the following two questions:
- Q2:
Is the algebra invariant under commuting moves in any other type?
- Q3:
Is the algebra invariant under other braid moves?
In section 3, we answer those questions applying the different techniques developed in section 2. In theorem 3.2, considering the geometric realization of a Coxeter system (assuming all the technical conditions described in [3]), we prove that the algebra is invariant under commuting moves, independently of the particular type of the system. Although we already prove this for type in [12], the proof is different in the general case. In theorem 3.2 we use the new techniques developed in this article. Finally in proposition 3.3 we answer We provide an explicit example and again, we use the techniques developed in section 2 to verify that the algebra is not always invariant under other braid moves.
1.2 Acknowledgment
This work was supported by FONDECYT de Postdoctorado 2024, N°3240046. ANID, Chile.
1.3 Notations and terminology
is a fixed field of characteristic different from
is the zero matrix.
We denote by the matrix given by
(1.3.1) -
A SLTM, is a strictly lower matrix, with entries in A SLTM, is a strictly lower matrix (of any size), with entries in We denote by , the set of all the SLTM, and by , the set of all the SLTM.
If is a SLTM, and then we denote by the matrix obtained from by erasing the rows and the columns of from the th to the th.
Given a SLTM, a SLTM, and a matrix we define the matrix by blocks, as follows:
1.4 Generalities
In this subsection we recall the main facts on the category introduced in [12].
Definition 1.1.
Given a SLTM, the nil graded algebra associated to is the commutative unitary algebra defined by generators and relations:
We denote this algebra by We define a grading on by for any
Note that definition 1.1 implies that the algebra is a nil graded algebra, that is, a graded algebra, where any homogeneous element with degree different from is a nilpotent element (see [12] for details). In particular, if then definition 1.1 implies that is generated by square nilpotent elements.
We denote by the family of all nil graded algebras defied by some SLTM, and the family of all nil graded algebras defined by some SLTM.
The following theorem was proved in [12]:
Theorem 1.2.
If is a SLTM, then the dimension of the algebra is and the set is a basis for
Definition 1.3.
Given two SLTM, a preserving degree homomorphism is any homomorphism of algebras such that for any If also is an isomorphism we say that is a preserving degree isomorphism.
Definition 1.4.
We define the category as the one whose set of objects is and whose morphisms are all the preserving degree homomorphisms between them. An isomorphism in is a preserving degree isomorphism between two members of
In the following we shall say that the algebras and are isomorphic if they are isomorphic as object of the category (that is, in the sense of definition 1.4).
In this article we are interested in to find out certain matrix criteria, to determine when two objects of are isomorphic. Given two SLTM, we denote by whenever the algebras are isomorphic. Relation is a equivalence relation on the set of all SLTM.
2 Isomorphism criteria
2.1 The fundamental equations
We shall assume that is a SLTM, is a SLTM, and are the generators defining and respectively. A preserving degree homomorphism defines a matrix by the equation
The following lemma provides a first matrix criterium to find out if a given matrix defines a homomorphism or not.
Lemma 2.1.
Given a matrix, then the assignment
defines a preserving degree homomorphisms of algebras if and only if
Proof.
The assignment defines a morphism in the category if and only if for each we have:
If we develop the left side of equation 2.1.3, we obtain:
Now by relation 1.4.1, we obtain
On the other hand, if we develop the expression on the right in equation 2.1.3, we obtain:
If we develop the right side of equation 2.1.4, we obtain
Again by relation 1.4.1, we obtain
or equivalently:
for any and any pair such that as desired. ∎
The following theorem is a direct consequence of lemma 2.1. In this case we shall assume that and are both SLTM.
Theorem 2.2.
Given a matrix, then the assignment
defines a preserving degree isomorphisms of algebras if and only if
and is an invertible matrix.
2.2 Small cases
Proposition 2.3.
There is only one class in under relation
Proof.
Let
Then the matrix
defines an isomorphism ∎
The case is a little bit more complicated. We divide the analysis in a series of lemmas:
Lemma 2.4.
There is at least two classes in under relation
Proof.
Note that the matrices
are not related by In fact if there was an invertible matrix satisfying equation 2.1.6, then we have for any
If we assume that then by the second and third equations in 2.2.1, we have and but this contradicts the first equation in 2.2.1, therefore for all implying that is not invertible. ∎
Given a SLTM, we denote
Lemma 2.5.
Given a SLTM, such that and then we have
Proof.
1.
If then and the matrix
defines an isomorphism
2.
If then the matrix
defines an isomorphism
Lemma 2.6.
Given a SLTM, such that or then
Proof.
The following list shows all the different cases for satisfying the hypothesis of this lemma, and the corresponding matrix that defines an isomorphism
(Note that in the last case, we have since ). ∎
Proposition 2.7.
There are exactly two classes in under relation
Proof.
The results provided in this section invite us to ask ourselves two natural question:
- Q1:
Is the number of classes in (under relation ) always finite?
We do not have an answer for those questions yet, but with the techniques developed we are able to find at least a lower bound for that number, see subsection 2.4 below.
2.3 The class of
In this section we describe a very explicit criterium (theorem 2.9) for determining when a given SLTM, is in the same class of the matrix This criterium is a generalization of lemma 2.6.
We first need the following definition:
Definition 2.8.
Given a SLTM, a scalar and indices we define the triangular subdeterminant
Note that if we take for we recover the number defined in subsection 2.2.
Let and the generators of and respectively.
Theorem 2.9.
Given a SLTM, then if and only if one of the following conditions holds
1.
for each pair of indices
2.
Given a SLTM, then for each pair of indices such that we have that for each
Lemma 2.10.
If is a SLTM, such that then one of the following conditions holds
1.
for each pair of indices
2.
For each pair of indices such that we have that for each
Proof.
If is an isomorphism, then by theorem 2.2 we have
for each and each Lets assume that there is a for some pair with Since the matrix is invertible, there is a such that therefore equation 2.3.1 implies that
On the other hand if we take in equation 2.3.1, we can also conclude that therefore
since equation 2.3.1 is true for each for each and each Now equation 2.3.2 implies that
Therefore we have
Definition 2.11.
Let be a SLTM.
1.
A leader of is a pair such that and for all We denote by the set of all leaders of
2.
We define the graph of leaders of as the directed graph where the set of vertices is and where arrows are defined by:
We also denote by if there is a finite sequence such that and for each
3.
Let a connected subgraph of We denote by the set given by:
If the set is not empty, we say that it is a chain of leaders of If is a (not trivial) connected component of then we say that is a maximal chain of leaders. If are two different connected components of then we say that and are two independent chains of leaders of
4.
A simple chain of leaders of is a subset say with and such that
5.
A ramified chain of leaders of is any chain of leaders that is not simple.
Lemma 2.12.
Given a SLTM, we have that:
1.
Relation satisfy the following properties:
(2.3.3) and
(2.3.4) -
2.
The following relation:
(2.3.5) is a partial order on the set of vertices of
3.
For each there is a unique minimal element of (relative to ), such that Moreover if are in the same connected component of then there is a unique minimal element of such that and
4.
If in not a minimal element of under then the set
is a simple chain of leaders of
5.
is a minimal element if and only if there is no such that and that means that the th row of is zero.
The proof of lemma 2.12 is direct, we left to the readers.
Example 2.13.
1.
Let
(2.3.6) then In this case is itself a ramified chain of leaders of The associated graph looks like this:
Note that is composed by two simple chains: and
2.
Let
(2.3.7) Then In this case the associated graph is disconnected and it looks like this:
The sets are two independent simple chains of leaders such that
3.
Let
(2.3.8) In this case then is itself a simple chain of leaders. The associated graph in this case looks like this
4.
Consider the matrix
(2.3.9) In this case is itself a ramified chain, composed by two simple chains where
The associated graph in this case looks like this:
5.
Consider the matrix:
(2.3.10) Then the set is a ramified chain of leaders of The directed graph associated is given by:
Note that if we write and then are two (not independent) simple chain of leaders of such that
6.
Consider the matrix:
(2.3.11) Then the set is composed by two independent chains of leaders. The directed graph associated is given by:
Note that is divided in two independent (maximal) chains:
Also note that is a simple chain, while is a ramified chain, with where
Lemma 2.14.
Let be a SLTM. If for each pair with we have that for each then for each pair with such that we have that for each
Proof.
Let and suppose that In particular this implies that there is a such that By our hypothesis, for each and this implies that if and only if In particular, we have that This implies that in In fact, if we have that Otherwise, if then since there is a such that and again this implies that By a recursive argument, we can find a finite sequence: such that for each that is in other words
1.
Fix an If we assume that in then we have that:
The last equations imply that:
Therefore we have:
2.
Fix an . If there is a finite sequence: such that for each that is, then we have:
(2.3.12) (2.3.13) (2.3.14) From the equations in 2.3.12 we obtain respectively:
(2.3.15) Using recursively equations 2.3.13 and 2.3.14 we obtain respectively:
(2.3.18) and
(2.3.19) Finally we obtain:
as desired.
Lemma 2.15.
Assuming that is itself a simple chain of leaders, say such that, for each we have for each Then
Proof.
First note that the hypothesis implies that the presentation of is given by the relations
also the hypothesis implies that the all the entries of are determined by the values of with Since is a chain we have that for each We also denote
We define for each minimal element (in particular for ), the element Being minimal means that there is no such that in particular by lemma 2.12, we know that
For we define the element recursively by:
We claim that the assignment defines an isomorphism By equation 2.3.20, it is enough to prove that for each
For note that and then
For we assume that then
One can check recursively (similarly to the proof of lemma 2.14) that:
that is
and so on.
Therefore
and
Now note that condition for each implies that
Moreover, we can use equation recursively to check that
for any
In fact, it is clear that
Assuming that equation 2.3.25 is true, for any then we have:
therefore
and equation 2.3.25 is true for any for any
Example 2.16.
Let be the matrix of equation 2.3.8, then then the presentation of is given by the equations
Proposition 2.17.
Let be a SLTM, and assume that for each with we have that for each then
Proof.
For each minimal element we define Again, being minimal means that there is no a such that therefore lemma 2.12 implies that
For not minimal, we define recursively:
By property 2.3.4 of lemma 2.12, is well defined by equation 2.3.26.
By item 4 of lemma 2.12, we know that the set is a simple chain of leaders, therefore the fact that follows analogously to the proof of this in lemma 2.15. Note that each independent chain can be treated independently. Therefore we have that the assignment defines an homomorphism
It is clear that the associated matrix is invertible, therefore is an isomorphism. ∎
Example 2.18.
Example 2.19.
Example 2.20.
Let be the matrix of equation 2.3.9. As we mentioned in example 2.13 is a ramified chain, composed by two (not independent) simple chains, and One can check that satisfies all the conditions of proposition 2.17. The isomorphism described in the proof of proposition 2.17, is given by:
The associated matrix is
Example 2.21.
Let the matrix of equation 2.3.10. As we have mentioned in example 2.13, is a ramified chain composed by the two (maximal) simple chains: and Note that in this case the minimal vertices are all of them but are isolated vertices of Also note that satisfies the hypothesis of proposition 2.17. Therefore
Following the proof of proposition 2.17, we have that:
Using the recursive definition provided in equation 2.3.26 in the proof of proposition 2.17, we have for :
Analogously for we obtain:
2.4 A lower bound for the number of classes
In this subsection we show the following:
Theorem 2.23.
There are at least classes in .
To do so, we shall show that is a set of representatives of pairwise different classes under relation
Lemma 2.24.
Considering the matrix we have that if and only if
Proof.
Is not difficult to see that the matrix satisfies the conditions of theorem 2.9 if and only if ∎
Proposition 2.25.
The matrices and are related under if and only if
Proof.
By lemma 2.24, we know that whenever Therefore we only have to prove that the matrices and for are not related by
If we assume that there is an isomorphism then by theorem 2.2, we have
In particular if we fix a then for each pair equation 2.4.1 reduces to
If we assume that there is a such that then equation 2.4.2 implies that whenever and On the other hand, if we assume that for all then necessarily since But if we replace in equation 2.4.1, we obtain and therefore for all but this is a contradiction with the definition of the matrix The last analysis implies that for each there is a such that and for each with
Now we replace in equation 2.4.1, and such that then we obtain Since by definition of the matrix then we conclude that That is valid for each Therefore, since we conclude that By the last analysis we also conclude that for each pair we have In particular, we conclude that there are exactly rows such that
If we replace in equation 2.4.1 we obtain
and by definition of the matrices and the indices we obtain
Note that by the symmetry of equation 2.4.3 we can omit the hypothesis that
Going back again to equation 2.4.1 but this time taking we obtain
By definition of the matrix and by the facts that for each and we obtain more precisely that
If we take for some we obtain
Let us fix such that (we know there is at least two rows satisfying that condition. Then we have:
- 1.
- 2.
The last two cases implies that implies that and then there are at most rows such that but this is a contradiction, since we know there are exactly rows satisfying that condition. By the last analysis we conclude that and are not isomorphic, and then ∎
2.5 Elementary triangular operations
In this section we define a series of matrix operations called Elementary triangular operations (ETO), that preserves the class under relation At the end of the subsection we prove for any pair that for having is necessary and sufficient that there exists of a (finite) sequence of ETO that transform into
Definition 2.26.
Let be a SLTM, fix an and an scalar We define the matrix as the SLTM, where
Example 2.27.
Consider the matrix
then
Lemma 2.28.
Let ba a SLTM, fix and an scalar Then
Proof.
Let Let and the generators of and respectively. We claim that the assignment
defines an isomorphism To see that recall that by equation 2.1.3, we only have to check that and the associated matrix is invertible. Now note that
1.
If then clearly
2.
If then we have
and
3.
if then we have
and
By the last analysis, we conclude that the assignment defines an homomorphism Is not difficult to see that the associated matrix is invertible, in fact Therefore we conclude that defines an isomorphism ∎
Definition 2.29.
Let be a SLTM, and fix If the matrix satisfies the following conditions:
1.
The entry for each
2.
The entry for each
then we define the matrix as the SLTM, where
Example 2.30.
Consider the matrix
and take and Is not difficult to see that satisfies the conditions of definition 2.29. Then the matrix is well defined. Moreover we have
Lemma 2.31.
Let be a SLTM, let If satisfies the conditions of definition 2.29, then
Proof.
Let Let and the generators of and respectively. We claim that the assignment
defines an isomorphism To see that, note that by our hypothesis we have:
1.
For we have
and
2.
For we have
and
3.
For it is immediate that
4.
For we have
and
5.
For we have
and
The last analysis implies that the assignment defines a homomorphism Is not difficult to see that the associated matrix is invertible, in fact Therefore defines an isomorphism ∎
Recall that for a given SLTM, we denote
Definition 2.32.
Let be a SLTM, take an scalar and assume that there are indices such that for each If one of the following conditions holds:
1.
2.
and for each
Then we define the matrix as the SLTM, given by
Example 2.33.
Consider the matrix
Let and Note that the conditions of definition 2.32 hold. In this case Therefore we have
Example 2.34.
Therefore we have
Lemma 2.35.
Let a SLTM, let be a pair of indices and an scalar. Assuming that the conditions of definition 2.32 hold, then we have that
Proof.
Let Let and the generators of and respectively. We divide the analysis in two cases:
1.
If then we have that the matrix is given by:
We claim that the assignment:
defines an isomorphism In fact, if then it is clear that
We also can check for that
and
Finally for we have
and
Therefore, the assignment defines a morphism Is not difficult to see that the matrix is invertible, moreover then is an isomorphism.
2.
If and for each then the matrix as the SLTM, is given by
We claim that the assignment:
defines an isomorphism In fact, if then it is clear that
We also can check for that
On the other hand
Note that by our hypothesis, we have then
then as desired.
Finally for we have
and
Therefore, the assignment defines a morphism Is not difficult to see that the matrix is invertible, moreover then is an isomorphism.
Definition 2.36.
We call Elementary triangular operations (ETO for short) any of the following matrix assignments (whenever they are well defined):
Given two SLTM, and we say that they are equivalent by ETO if there is a finite sequence of ETO that transform into We denote by whenever and are equivalent by ETO.
Example 2.37.
Consider the matrices
then In fact one can check that
Definition 2.38.
In the following we shall denote respectively by the matrices given by:
Note that the matrix described in the proof of lemma 2.28 is of the form the matrix described in the proof of lemma 2.31 is of the form and the matrix described in the proof of lemma 2.35 is of the form
Theorem 2.39.
Let be two SLTM, then if and only if
Proof.
By lemmas 2.28, 2.31 and 2.35 we have that if then Moreover lemmas 2.28, 2.31 and 2.35 implies that for two SLTM, we have that if and only if there is an invertible matrix that can be written as a (finite) product of matrices of the form and
If then there exist an invertible matrix satisfying equation 2.1.6. Since is invertible, it can be factored as a product of elementary matrices (in the usual sense), Now the result follows from the fact that the matrices and are elementary matrices and any other elementary matrix can be written as a product of matrices of the form and ∎
3 Applications To Soergel Calculus
Fix a Coxeter system of finite rank (with for certain set of indices ), and a field where the number has sense in for each pair Following [6], the geometric realization of over a is a triple composed by a vector space subsets and where:
and the equations:
define representations of the group on and respectively. For any and any let
We are avoiding a technical definition on the diagrammatic Soergel category since it is well known. We invite the readers to look for more information in the original work of Elias and Williamson [3]. In this section we shall assume that the ground field has characteristic different from and satisfies all the technicalities described in [3]. Let the free monoid generated by the symbols In [12], for each element we define the SLTM:
and the corresponding nil graded algebra We have proved in [12], that the algebra corresponds to the Gelfand-Tsetlin subalgebra of the endomorphism algebra in the diagrammatic Soergel Category of Elias and Williamson (see [3]), relative to the Libedynski’s light leaves basis (see [8]) and the set of Jucys-Murphy elements defined by Ryom-Hansen in [19].
In [12] we show that for the case of a Coxeter system of type the algebras are invariant under commuting moves, and we proved that by defining explicit isomorphisms. Now in theorem 3.2 we generalize this fact for any other type. In this case we use the techniques developed in this article, to obtain a cleaner proof.
Let us recall first, certain technical definitions coming from [12]:
Definition 3.1.
1.
Let we define the column vector associated to as the matrix given by:
2.
Given two elements of say and we define the matrix by
where
In [12], we show that for any composed expression we have that:
Moreover in [12], we proved that for we have the following (kind of) associative property:
where
Now we are ready to prove the following:
Theorem 3.2.
For any Coxeter system the algebra is invariant under commuting moves.
Proof.
Let such that that its, the elements and of the group commute. Let and then are elements of that differ by a commuting move.
1.
Let and Then and Therefore:
(3.0.3) where
2.
Note that by definition and Therefore we have:
By the same reason, we have that two first rows of are given by:
analogously the two first rows of are given by:
3.
Note that for every we have that Let us denote
then
4.
Following with the notation of the s, we have that the rows of the matrices from the third to the last one are given respectively by:
and
But since and commute, we clearly have that
- 5.
Theorem 3.2 naturally opens the following question:
- Q3:
Is the algebra invariant under other braid moves?
Fortunately we have an answer for Q3, although it is not a positive one:
Proposition 3.3.
The algebra is not invariant under braid moves.
Proof.
It is enough to provide an example where two equivalent by braid moves expressions and generate two not isomorphic algebras and
Let be the Coxeter system of type Let and then:
1.
It is clear that and are equivalent by braid moves.
2.
The corresponding matrices and are given by:
3.
Considering the matrices defined by:
one can check that they are well defined (see subsection for details) and by definition More specifically we have:
4.
Let be the entries of let be the entries of the matrix By theorem 2.39, if the algebras and were isomorphic, then necessarily there will be an isomorphism Assuming that, then the associated matrix would satisfy equation 2.1.6 of theorem 2.2. In particular we would have for and any pair that:
(3.0.4) -
5.
Equation 3.0.4 in particular implies that:
- 6.
7.
Let and be the sets of generators of the algebras and respectively. From the last two items, we conclude that if there exists an isomorphism then necessarily for In particular this implies that the matrices and are in the same class under relation but this contradicts theorem 2.9. In fact
The analysis above implies that and are not isomorphic. ∎

### v2

@@
 Abstract
-The category was introduced in [12] in order to provide a structural setting to the study of the many Gelfand-Tsetlin subalgebras appearing in the context of the diagrammatic Soergel category of Elias and Williamson [3].
-The category have as objects, all that we called nil graded algebras associated to a triangular matrix and as morphisms, all the preserving degree homomorphisms of graded algebras between them.
-In this article we develop a series of isomorphism criteria on that we used later to find out relevant information on the Gelfand-Tsetlin subalgebras of the diagrammatic Soergel category.
+The category was introduced in [15] in order to provide a structural setting to the study of certain commutative algebras appearing in the context of the diagrammatic Soergel category of Elias and Williamson [3].
+The category have as objects, all that we called nil graded algebras associated to triangular matrices and as morphisms, all the preserving degree homomorphisms of graded algebras between them.
+In this article we develop a series of isomorphism criteria on and we show some of the applications relative to the search of a complete classification.
+keywords: Commutative graded algebras, Jucys-Murphy elements, Soergel Calculus.
 Contents
@@
 1.1 Motivation
-The diagrammatic Soergel Category was defined in 2016 by Elias and Williamson [3] in order to provide a diagrammatic approach to the study of the category of Soergel bimodules, a well known incarnation of the Hecke category, [18].
+The diagrammatic Soergel Category is a linear and monoidal category, it was defined in 2016 by Elias and Williamson [3] in order to provide a diagrammatic approach to the study of the category of Soergel bimodules, a well-known incarnation of the Hecke category, [21].
 Since then, the category have gained attention of both algebraists and algebraic geometrist, due to its intrinsic connection with both branches of study.
+For further studies on this topic, we recommend to see [4] and [12] for example.
 A particular interest is the study of the internal structure of the category itself, that is, the structure of the many Hom spaces and more particularly of the many End algebras.
 In this context, for example, the diagrammatic setting provided by Elias and Williamson have been crucial to find out and prove the cellularity of the Light leaves bases of Libedinsky [9],[3].
-Another example of the power of this diagrammatic approach is the discovery due to Ryom-Hansen, of explicit Jucys-Murphy elements relative to the Libedinsky’s bases [19].
-In [10], Libedinsky and Plaza, found out a hard connection between the category and what they called the Blob category, another diagrammatic category, inspired in the well known KLR-algebras of Khovanov, Lauda and Rouquier.
-The similar patterns that they observed in both categories motivate them to define the Blob vs Soergel conjecture, where roughly speaking, they declare that there is an explicit incarnation of the category (in type ), in terms of KLR-diagrams.
-Some of the first evidences on the veracity of the aforementioned conjecture can be found in the work of Espinoza and Plaza [4], where explicit isomorphisms were built between the corresponding Gelfand-Tsetlin subalgebras, that is, the algebras generated by the Jucys-Murphy elements, coming on one side from Two-color Soergel Calculus and on the other of certain idempotent truncation of the KLR version of the blob algebra (see [17]).
-In [13], Plaza, Ryom-Hansen and the author, obtain isomorphisms in a higher level, this time between certain endomorphism algebras of in type and the corresponding idempotent truncation of blob algebras.
-In both cases the procedure implies the search of optimal presentations in terms of generators and relations for the algebras involved, therefore, in both [4] and [13], the reader will find, not only a clear proof of each isomorphism, but also deep information on the structure of the corresponding algebras.
-Recently in [1], Bowman, Cox and Hazi have proved the Blob vs Soergel conjecture, they in fact, proved a more general result, obtaining as a consequence the equivalences predicted by Libedinsky and Plaza.
-The smart procedure used in [1], based in alcove geometry and paths configurations, allows the authors to cover all the cases in just one go, while in [4] and [13] only particular cases were covered.
+Another example of the power of this diagrammatic approach is the discovery due to Ryom-Hansen, of explicit Jucys-Murphy elements relative to the Libedinskyβs bases [22] (In the following when we mention Jucys-Murphy elements in we refer to those developed by Ryom-Hansen in [22]).
+Finally we would like to mention, probably the most important application of this diagrammatic setting, the disproof of the famous Lusztigβs conjecture, given by Williamson in [24].
+In [13], Libedinsky and Plaza, found out a connection between the category and what they called the Blob category, another diagrammatic category, inspired in the well-known KLR-algebras of Khovanov, Lauda and Rouquier.
+The similar patterns that they observed in both categories motivate them to define the Blob vs Soergel conjecture, where roughly speaking, they declare that these two categories are equivalent.
+Some of the first evidences on the veracity of the aforementioned conjecture can be found in the work of Espinoza and Plaza [5], where explicit isomorphisms were built between the subalgebras generated by the Jucys-Murphy elements, coming on one side from Two-color Soergel Calculus and on the other of certain idempotent truncation of the KLR version of the blob algebra (see [20]).
+In [16], Plaza, Ryom-Hansen and the author, obtain isomorphisms in a higher level, this time between endomorphism algebras of in type and the corresponding idempotent truncation of blob algebras.
+In both cases the procedure implies the search of adequate presentations in terms of generators and relations for the algebras involved, therefore, in both [5] and [16], the reader will find, not only a clear proof of each isomorphism, but also more information on the structure of the corresponding algebras.
+Some years ago, Bowman, Cox and Hazi have proved in [1], the Blob vs Soergel conjecture, they in fact, proved a more general result, obtaining as a consequence the equivalences predicted by Libedinsky and Plaza.
+The procedure used in [1], based in alcove geometry and paths configurations, allows the authors to cover all the cases in just one go, while in [5] and [16] only particular cases were covered.
 On the other hand the general approach developed in [1] do not implies the aforementioned internal information on the corresponding algebras obtained in [5] and [16] or any generalization of that.
-Therefore, we consider as an open problem the search of this internal information for all the cases no covered in [4] and [13], that is in principle, to find out optimal presentations for all the algebras related with the Blob vs Soergel conjecture.
-In the last few years, the author have been developing certain strategies to address this problem.
-For example in [11], presentations for the Gelfand-Tsetlin subalgebras of a family of idempotent truncations of the generalized blob algebras were found, later in [12], working on the side of the category the author developed a series of techniques and procedures to obtain optimal presentation for the Gelfand-Tsetlin subalgebras of the many End algebras for any type (far beyond the initial objective: type ).
-Both [11] and [12] are natural generalizations of [4] and particularly they seem to imply a hard connection between the light leaves bases of Libedinsky and the graded cellular bases built by Ryom-Hansen and the author in [14], just as was previously observed in [4] and [13], in that case, with respect to the graded cellular bases developed by Plaza and Ryom-Hansen in [17].
-The category was introduced in [12] in order to provide a structural setting to the study of the many Gelfand-Tsetlin subalgebras appearing in the context of the diagrammatic Soergel category The category have as objects, all that we called nil graded algebras associated to a triangular matrix and as morphisms, all the preserving degree homomorphisms of graded algebras between them.
-Naturally, the aforementioned Gelfand-Tsetlin subalgebras of are (the main) objects in our category, and all the knowledge of developed in [12] allowed us to find out the desired presentations for those algebras.
-On the other hand, the category is by itself an interesting object of study and the results obtained in [12] open the door to a deeper analysis on it.
-Being more specific, the objects of the category are commutative graded algebras defined by a presentation with a finite set of generators and certain quadratic relations codified by a strictly triangular matrix.
-In [12] we define and make a first study of those algebras, obtaining a series of techniques and corresponding applications to the study of the many Gelfand-Tsetlin subalgebras of the category Since each objects of is totally determined by a codifying triangular matrix, it seem plausible to ask when two of those objects are isomorphic and which role play the corresponding codifying matrices in this aspect.
-To answer this question is the main motivation of this and other works we have in progress.
-In section 2, we develop a series of isomorphism criteria on that partially answer the above question.
-In Theorem 2.2 we obtain an absolute (but not easy to apply) matrix criterion to determine if two objects are or not isomorphic, this is the key result that allowed us to find out different approaches to a complete classification of the objects of Applying conveniently theorem 2.2, we could describe completely the class of the algebra associated to the zero matrix obtaining a very clear and easy to apply criterium (Theorem 2.9) to determine if the algebra is isomorphic to or not, depending on the entries of the codifying matrix The case of the matrix although very particular, can be cleverly used to determine if other two objects of are or not isomorphic (see for example the proof of proposition 3.3).
-Another fundamental result of this article is Theorem 2.39, where we describe another absolute criterium (again, not always easy to apply) to determine if two objects are or not isomorphic.
-In the case of theorem 2.39, the criterion states that two objects and are isomorphic if and only if, there is a finite sequence of certain matrix operations (that we called elementary triangular operations), that progressively transform the matrix into the matrix
-Our first examples are developed in subsection 2.2, where the smaller cases were studied.
-The results that we found there inspire all the analysis we did in this article.
-In particular, they open interesting questions for further works, for example, since we observe that for small matrices, we have a very small number of isomorphic classes in then is natural to ask:
+Therefore, we consider as an open problem the search of this internal information for all the cases no covered in [5] and [16], that is, in principle, to find out presentations for all the algebras related with the Blob vs Soergel conjecture.
+In the last few years, we have been developing certain strategies to address this problem.
+For example in [14], we obtained presentations for the subalgebras generated by Jucys-Murphy elements in certain family of idempotent truncations of the generalized blob algebras.
+Later in [15], working on the side of the category we developed a series of techniques and procedures to obtain presentations for the subalgebras generated by Jucys-Murphy elements (from now the algebras), in the many End algebras for any type (far beyond the initial objective: type ).
+Both [14] and [15] are natural generalizations of [5].
+In general the subalgebra generated by the Jucys-Murphy elements in a cellular algebra is a commutative subalgebra that under certain conditions, it could be a maximal commutative subalgebra of (see [18]).
+One can expect that subalgebra contains important information on the algebra itself.
+On the other hand Jucys-Murphy elements are very relevant in the study of the Representation Theory of the corresponding cellular algebra (see [18]).
+We think that all the study that we are doing on algebras in the context of the category will help to learn more on this important category.
+The category was introduced in [15] in order to provide a structural setting to the study of the many algebras appearing in the context of the diagrammatic Soergel category The category have as objects, all that we called nil graded algebras associated to triangular matrices and as morphisms, all the preserving degree homomorphisms of graded algebras between them.
+Naturally, the aforementioned algebras of are (the main) objects in our category, and all the knowledge of developed in [15] allowed us to find out the desired presentations for those algebras.
+In particular we would like to use our work in the search of a complete classification of the many algebras of the Soergel category
+On the other hand, the category is, in our opinion, by itself an interesting object of study, since their objects are basically certain quotients of polynomial algebras on several variables, they have potential connections with other branches of study, for example with Algebraic Geometry.
+To obtain tools to classify the objects on could help, for example, to obtain new tools for analysis of certain types of Algebraic Varieties.
+Being more specific, the objects of the category are commutative graded algebras defined by a presentation with a finite set of generators and certain quadratic relations codified by a strictly lower triangular matrix.
+The dependence on the codifying matrices, invite us to ask how those matrices, could be relevant when we want to know if two of those algebras are isomorphic or not.
+Moreover, we would like to know if a complete classification of the objects of is possible, and which role play the corresponding codifying matrices in these aspects.
+In this article we give some first steps in the search of answers for those questions.
+The content of this article is organized as follows:
+- β’
+β’
+In subsection 2.2, we obtain our first isomorphism criterion.
+In Theorem 2.6 we provide an absolute (but not easy to apply) matrix criterion to determine if two algebras in are or not isomorphic.
+Theorem 2.6, basically asserts that two algebras and in the category are isomorphic if and only if there exist an invertible matrix whose entries satisfy certain system of quadratic equations, determined by the entries of and
+β’
+In subsection 2.3, we define a set of matrix operations, that we called Elementary triangular operations (or ETOβs for shortness).
+This matrix operations act on the set of triangular matrices, each one subject to certain restrictions.
+In Definition 2.17, we define Equivalence by ETOβs, an equivalence relation between triangular matrices.
+In Theorem 2.19, we show that if two matrices are equivalent by ETOβs, then the corresponding algebras are isomorphic as objects of the category
+β’
+In Proposition 2.20, we connect the matrix coming from Theorem 2.6 and ETOβs.
+We prove that if is any elementary matrix, defining an isomorphism then it implies that the matrices and are equivalent by ETOβs.
+Since any isomorphism is determined by an invertible matrix (Theorem 2.6), and any invertible matrix can be factored as a product of elementary matrices, it seem sensible to think that the existence of an isomorphism between and will imply that the matrices are necessarily equivalent by ETOβs.
+We declare this in Conjecture 2.21.
+We do not have a proof of this yet, but we have an important set of evidences on its veracity, we show some of them in this article.
+β’
+In Section 3 we show some applications of the criteria obtained in Section 2.
+In Subsection 3.1, we obtain a complete classification of algebras where is a lower matrix of size or (see Propositions 3.1 and 3.3).
+We also prove for those cases the veracity of Conjecture 2.21 (see Corollaries 3.2 and 3.4).
+β’
+In Subsection 3.2, we use Theorem 2.6 to deduce an absolute criterion, depending explicitly on the entries of a matrix to determine if the algebras and are isomorphic or not.
+Theorem 3.12, is a clear example of what we were talking about, when we propose ourselves to look for criteria coming from the codifying matrices, to determine when two algebras in are isomorphic or not.
+Although it only cover one isomorphism class, we consider this result as a challenging invitation to obtain such as explicit criteria for other isomorphism classes.
+This is the motivation to keep on working in the search of a complete classification of the objects in On the other hand Proposition 3.14 provide another evidence on the veracity of Conjecture 2.21.
+In this case, Proposition 3.14 asserts that if an algebra is isomorphic to then necessarily the matrices and are equivalent by ETOβs.
+β’
+As we mentioned above, in Subsection 3.1, we obtained a complete classification for algebras codified by matrices of size or In fact, Propositions 3.1 and 3.3, provide explicitly the number of classes for those sizes.
+In both cases, the number is finite and it is independent on the ground field.
+Those facts invite us to open the following questions:
 - Q1:
 Is the number of classes always finite?
-Although we do not have an answer for this question yet, in subsection 2.4, we obtain a lower bound for the number of classes depending on the size of the codifying matrices.
-In theorem 2.23, we describe an explicit set of representatives of different classes, for each size We left Q1 open, together with the search of a complete set of representatives and explicit criteria to describe any isomorphic class such as we did in the case of the matrix
-On the other hand, in section 3, we advocate our attention to apply the results developed in section 2, to the study of the Gelfand-Tsetlin subalgebras coming from the diagrammatic Soergel category In [12], for each element we a define the SLTM, and the corresponding nil graded algebra We have proved in [12], that the algebra corresponds to the Gelfand-Tsetlin subalgebra of the endomorphism algebra in the diagrammatic Soergel Category of Elias and Williamson, relative to the Libedynski’s light leaves basis and the set of Jucys-Murphy elements defined by Ryom-Hansen (See [3],[8] and [19]).
-In [12] we show that for the case of a Coxeter system of type the algebras are invariant under commuting moves. That result naturally open the following two questions:
 - Q2:
-Is the algebra invariant under commuting moves in any other type?
-- Q3:
-Is the algebra invariant under other braid moves?
-In section 3, we answer those questions applying the different techniques developed in section 2.
-In theorem 3.2, considering the geometric realization of a Coxeter system (assuming all the technical conditions described in [3]), we prove that the algebra is invariant under commuting moves, independently of the particular type of the system.
-Although we already prove this for type in [12], the proof is different in the general case.
-In theorem 3.2 we use the new techniques developed in this article.
-Finally in proposition 3.3 we answer We provide an explicit example and again, we use the techniques developed in section 2 to verify that the algebra is not always invariant under other braid moves.
-1.2 Acknowledgment
-This work was supported by FONDECYT de Postdoctorado 2024, N°3240046. ANID, Chile.
+Is the number always independent of the ground field?
+We do not have answer for those questions yet, but we expect to address this problem in future works.
+For instance, motivated uniquely by our intuition, we define Conjecture 3.17, where we state that the number of isomorphism classes for depends on the ground field.
+Moreover the number is finite if and only if the ground field is finite.
+β’
+Independently on if the number is finite or not, in Subsection 3.3, we obtain a lower bound for the number of classes for each size that is independent on the ground field.
+In Theorem 3.16, we describe an explicit set of representatives of different classes.
+This result will be useful, when we have answered Q1 and Q2, and naturally look for the exact value of
+The main results of this article are:
 1.2 Notations and terminology
-is a fixed field of characteristic different from
-is the zero matrix.
+The following notation will be used along this article:
+β’
+is a fixed field of characteristic different from We also denote
+β’
+is the zero matrix and is the identity matrix, both of size
+β’
 We denote by the matrix given by
 (1.2.1) -
+β’
 A SLTM, is a strictly lower matrix, with entries in A SLTM, is a strictly lower matrix (of any size), with entries in We denote by , the set of all the SLTM, and by , the set of all the SLTM.
+β’
 If is a SLTM, and then we denote by the matrix obtained from by erasing the rows and the columns of from the th to the th.
-Given a SLTM, a SLTM, and a matrix we define the matrix by blocks, as follows:
+β’
+Given a SLTM, an scalar and indices we denote by the element When there is no possible confusion we only write instead of
+1.3 Acknowledgement
+It is a pleasure to thank Steen Ryom-Hansen for stimulating conversations on Jucys-Murphy elements and algebras.
+I would specially want to thank the anonymous referee for communicating detailed suggestions and remarks that helped to improve the clarity and presentation of this article.
+This work was supported by:
+β’
+FONDECYT de Postdoctorado 2024, NΒ°3240046, ANID, Chile.
+β’
+SubvenciΓ³n a la InstalaciΓ³n en la Academia 2024, NΒ°85240053, ANID, Chile.
+2 Isomorphism criteria
 2.1 Generalities
@@
 Relation is a equivalence relation on the set of all SLTM.
-2 Isomorphism criteria
-2.1 The fundamental equations
-We shall assume that is a SLTM, is a SLTM, and are the generators defining and respectively. A preserving degree homomorphism defines a matrix by the equation
-The following lemma provides a first matrix criterium to find out if a given matrix defines a homomorphism or not.
+2.2 Fundamental equation
+We shall assume that is a SLTM, is a SLTM, and are the generators defining and respectively.
+Note that any morphism defines a matrix by the equation
+In particular any morphism is totally determined by a linear transformation On the other hand, a given linear transformation not necessarily defines a morphism in the category The following lemma provides a first matrix criterion to find out if a given matrix defines a morphism via equation 2.2.1, or not:
 Lemma 2.5.
@@
 If we develop the left side of equation 2.2.3, we obtain:
-Now by relation 1.4.1, we obtain
+Now by relation 2.1.1, we obtain
 On the other hand, if we develop the expression on the right in equation 2.2.3, we obtain:
 If we develop the right side of equation 2.2.4, we obtain
-Again by relation 1.4.1, we obtain
+Again by relation 2.1.1, we obtain
 or equivalently:
-for any and any pair such that as desired. ∎
-The following theorem is a direct consequence of lemma 2.1. In this case we shall assume that and are both SLTM.
+for any and any pair such that as desired. β
 Theorem 2.6.
-Given a matrix, then the assignment
-defines a preserving degree isomorphisms of algebras if and only if
-and is an invertible matrix.
+Let be two SLTM. Given a matrix, then the assignment
+defines an isomorphisms if and only if is an invertible matrix whose entries satisfy the following system of equations:
+Proof.
+If is an isomorphism, then clearly the associated linear transformation (obtained by restriction) is invertible, and therefore the matrix (defined by equation 2.2.1) is invertible.
+Lemma 2.5 implies that the entries of satisfy equation 2.2.6.
+On the other hand if is an invertible matrix, and satisfies equation 2.2.6, then by Lemma 2.5, it defines a morphism Let be the inverse matrix of Let us define the function as follows
+where the sum runs over all the with is any scalar and is given by the equation
+Is not difficult to see that is the inverse function of Since is in particular an homomorphism of algebras, we conclude directly that is also an homomorphism of algebras.
+Since preserves degrees, so does therefore is a morphism in the category and this implies that is an isomorphism in the category β
+2.3 Elementary triangular operations
+In this section we define three matrix operations, that we call elementary triangular operations (ETOβs).
+Basically each ETO transform a SLTM, , into a new SLTM, As we will see in Definitions 2.7, 2.10 and 2.13, each ETO has certain restrictions, that is, we only are allow to apply them on matrices that satisfies a series of conditions.
+Anyway, we shall prove that, whenever this transformation is well-defined, we obtain a matrix such that Moreover we shall prove that if there is a finite sequence of, well-defined, ETOβs that transform into then we have that
+Definition 2.7.
+Let be a SLTM, fix an and an scalar We define the matrix as the SLTM, whose entries are given by:
+Example 2.8.
+Consider the matrix
+then
+Lemma 2.9.
+Let be a SLTM, fix and an scalar Then
+Proof.
+Let (as in definition 2.7). Let and the generators of and respectively. We claim that the assignment
+defines an isomorphism To see that, we only have to check that
+for each and that the associated matrix is invertible. Now note that
+- 1.
+- 2.
+- 3.
+By the last analysis, we conclude that the assignment defines an homomorphism Is not difficult to see that the associated matrix is invertible, in fact, the entries of are given by
+that is, then and we conclude that defines an isomorphism β
+In the following, we shall denote by the associated matrix of the isomorphism defined in equation 2.3.1. For example if and we have:
+Note that is an elementary matrix.
+Definition 2.10.
+Let be a SLTM, and fix If the matrix satisfies the following conditions:
+1.
+The entry for each
+2.
+The entry for each
+then we define the matrix as the SLTM, whose entries are given by:
+Example 2.11.
+Consider the matrix
+and take and Is not difficult to see that satisfies the conditions of Definition 2.10. Then the matrix is well-defined. Moreover we have
+Lemma 2.12.
+Let be a SLTM, let If satisfies the conditions of Definition 2.10, then
+Proof.
+Let (as in definition 2.10). Let and the generators of and respectively. We claim that the assignment
+defines an isomorphism To see that, we only have to check that
+for each and that the associated matrix is invertible. Now note that
+- 1.
+- 2.
+- 3.
+4.
+For we have on one side
+- 5.
+The last analysis implies that the assignment defines a homomorphism Is not difficult to see that the associated matrix is invertible, in fact, the entries of are given by
+Therefore and we conclude that defines an isomorphism β
+In the following we shall denote by the associated matrix of the isomorphism defined in equation 2.3.7. For example if and we have:
+Note that is an elementary matrix.
+Recall that for a given SLTM, we denote (see subsection 1.2).
+Definition 2.13.
+Let be a SLTM, take an scalar and assume that there are indices such that for each If one of the following conditions holds:
+1.
+2.
+and for each
+Then we define the matrix as the SLTM, whose engtries are given by:
+Example 2.14.
+Consider the matrix
+Let and Note that the conditions of Definition 2.13 hold. In this case Therefore we have
+Example 2.15.
+Therefore we have
+Lemma 2.16.
+Let a SLTM, let be a pair of indices and an scalar. Assuming that the conditions of Definition 2.13 hold, then we have that
+Proof.
+We shall assume that The case is easier and it is left to the readers.
+Assuming that and for each then the matrix is well-defined and its entries are given by equation 2.3.11 (see Definition 2.13). Let and the generators of and respectively. We claim that the assignment:
+defines an isomorphism As in Lemmas 2.9 and 2.12, we only have to check that
+for each and that the associated matrix is invertible. Now note that:
+- 1.
+2.
+If we have on one side:
+The last step is because equation 2.3.11, from where we know that for all for and Also by the restrictions given in definition 2.13, we have that for each
+On the other side, equation 2.3.12 implies that
+Note that by our hypothesis, we have then we can reduce the expression on the parenthesis in the first sum to obtaining:
+then equation 2.3.13 holds in this case.
+3.
+If we have on one side:
+since by equation 2.3.11, that for and
+By the above analysis we conclude that, the assignment defines an homomorphism Is not difficult to see that the matrix is invertible, in fact, the entries of are given by:
+therefore and we conclude that defines an isomorphism. β
+In the following we shall denote by the associated matrix of the isomorphism defined in equation 2.3.7. For example if and we have:
+Note that is an elementary matrix.
+Definition 2.17.
+We call Elementary triangular operations (ETO for short) any of the following matrix assignments (whenever they are well-defined):
+Given two SLTM, and we say that they are equivalent by ETO if there is a finite sequence of ETOβs that transform (progressively) into We denote by whenever and are equivalent by ETO.
+It is clear that defines an equivalence relation on the set of all SLTM. In particular note that
+Each of the assignments in equation 2.3.17 is well-defined if and only if the corresponding assignment in equation 2.3.16 is well-defined.
+Example 2.18.
+Consider the matrices
+then In fact one can check that
+Theorem 2.19.
+Let be two SLTM. If then
+Proof.
+If we assume that that is, there is a finite sequence of ETOβs, that transform into say where each is one of the three ETOβs described in Definition 2.17, and the corresponding matrix is well-defined, that is, respects the corresponding restrictions of the operation Then Lemmas 2.9, 2.12 and 2.16 imply that at each step The result follows by transitivity of relation
+β
+Proposition 2.20.
+Let be two SLTM. Assuming that there is an isomorphism such that, the associated matrix is an elementary matrix, then
+Proof.
+We fix a SLTM, T, and we left the matrix variable. Let be an elementary matrix. We separate our analysis in the following cases:
+- Case 1
+If for some and some (for notation, see just below of the proof of Lemma 2.9).
+If we assume that defines an isomorphism then the entries of must satisfy equation 2.2.6, where are replaced by the entries of our fixed matrix and are replaced by the entries of which are given explicitly in equation 2.3.4.
+Using equation 2.2.6 in this way, we have:
+- 1.
+- 2.
+3.
+For any other replacement of a triple such that we obtain equations implying that or equations of the form
+The above analysis implies that the entries of are given by equation 2.3.1, and then Therefore in this case.
+- Case 2
+If for some (for notation, see just below of the proof of Lemma 2.12).
+We apply a similar analysis as in Case 1, we assume that defines an isomorphism then the entries of must satisfy equation 2.2.6, where are replaced by the entries of our fixed matrix and are replaced by the entries of which are given explicitly in equation 2.3.9.
+Using equation 2.2.6 in this way, we have:
+- 1.
+- 2.
+- 3.
+- 4.
+- 5.
+- 6.
+7.
+If we replace a triple such that and in equation 2.2.6, we obtain
+(2.3.26) -
+8.
+If we replace a triple such that and in equation 2.2.6, we obtain
+(2.3.27) -
+9.
+If we replace a triple such that and and in equation 2.2.6, we obtain
+(2.3.28) -
+10.
+If we replace a triple such that and and in equation 2.2.6, we obtain
+(2.3.29) -
+11.
+For any other replacement of a triple such that we obtain equations implying that or equations of the form
+Note that equations 2.3.20, 2.3.21 and 2.3.22 imply that if the matrix defines an isomorphism then necessarily, the matrix must satisfy all the restrictions described in Definition 2.10.
+The above analysis imply that and therefore
+- Case 3
+If (for notation see just below of the proof of Lemma 2.16).
+The analysis is similar to those of Case 1 and Case 2, we have to replace the entries (defined in equation 2.3.14), in equation 2.2.6, collecting restrictions on and defining equations for the entries of We shall assume that the case is easier and it is left to the readers.
+In this case we have the following situations:
+- 1.
+2.
+If we replace a triple such that and in equation 2.2.6, we obtain
+(2.3.31) - 3.
+- 4.
+5.
+If we replace a triple such that in equation 2.2.6, we obtain
+(2.3.35) -
+6.
+For any other replacement of a triple such that we obtain equations implying that or equations of the form
+Note that equations 2.3.30, and 2.3.33 imply that if the matrix defines an isomorphism then necessarily, the matrix must satisfy all the restrictions described in Definition 2.10.
+The above analysis imply that and therefore
+- Case 4
+If then necessarily that is, there is a pair and an scalar such that
+(2.3.36) Working exactly as we did in the three previous cases, we obtain that:
+1.
+The entries of have the following restrictions:
+(2.3.37) In particular, they imply that for all and therefore the matrix is well-defined.
+2.
+The entries of depends on as follows:
+Now note that the entries of the matrix are given by:
+(2.3.39) In particular we have that for all and for all This implies that the matrix is well-defined.
+Now note that the entries of are given by:
+(2.3.40) Let us define the matrix Since the operation have no restrictions, is well-defined and its entries are given by:
+(2.3.41) Note in particular that for all therefore we can define the matrix whose entries are given by:
+(2.3.42) Finally we define the matrix By contruction and one can check that the entries of are given by
+(2.3.43) Comparing equations 2.3.39 and 2.3.43, one conclude that and therefore as desired.
+1.
+β
+Note that in Case 4 of the proof of Proposition 2.20, the matrix can be factored as follows:
+and that factorization yields the following well-defined process of transformations:
+That is, the restrictions and relations for matrices obtained, when the entries of are replaced in the system of equations 2.2.6, are coherent with the process of transformations induced by the factorization given in equation 2.3.44.
+More generally, any invertible matrix can be factored as a product of elementary matrices, and each elementary matrix can be written as product of matrices of the forms or Now if we assume that is an invertible matrix, that define an isomorphism for certain variable matrices then the system of equations 2.2.6, will imply certain restrictions and relations on and Based in the proof of Proposition 2.20, it is sensible to think that, there will be a factorization with matrices of one of the forms or yielding a well-defined process of transformations:
+At this point we do not have a well proof of that fact, but in an enormous set of explicit examples, we always obtain evidences of its veracity. Based on that, we define the following:
+Conjecture 2.21.
+(Isomorphism vs Equivalence by ETOβs) Let be two SLTM. If then
+3 Applications
+In this section we apply the isomorphism criteria obtained in Section 2, to give some first steps in the search of a complete classification of the objects of the category
 3.1 Small cases
+In the smallest case, that is, we can easily count exactly the number of classes:
 Proposition 3.1.
@@
 Then the matrix
-defines an isomorphism ∎
-The case is a little bit more complicated. We divide the analysis in a series of lemmas:
-Lemma 2.4.
-There is at least two classes in under relation
+defines an isomorphism β
+Corollary 3.2.
+If are any two matrices in If then
 Proof.
-Note that the matrices
-are not related by In fact if there was an invertible matrix satisfying equation 2.1.6, then we have for any
-If we assume that then by the second and third equations in 2.2.1, we have and but this contradicts the first equation in 2.2.1, therefore for all implying that is not invertible. ∎
-Given a SLTM, we denote
-Lemma 2.5.
-Given a SLTM, such that and then we have
-Proof.
-1.
-If then and the matrix
-defines an isomorphism
-2.
-If then the matrix
-defines an isomorphism
-Lemma 2.6.
-Given a SLTM, such that or then
-Proof.
-The following list shows all the different cases for satisfying the hypothesis of this lemma, and the corresponding matrix that defines an isomorphism
-(Note that in the last case, we have since ). ∎
+The case is a little bit more complicated, but again, we can count exactly the number of classes. For the rest of this subsection for a given SLTM, we denote (for notations, see subsection 1.2).
 Proposition 3.3.
@@
 Proof.
-The results provided in this section invite us to ask ourselves two natural question:
-- Q1:
-Is the number of classes in (under relation ) always finite?
-We do not have an answer for those questions yet, but with the techniques developed we are able to find at least a lower bound for that number, see subsection 2.4 below.
+We divide the proof in a series of steps:
+- Step 1
+Note that the matrices
+are not related by In fact if there was an invertible matrix satisfying equation 2.2.6, then we have for any
+(3.1.3) If we assume that then by the second and third equations in 3.1.3, we have and but this contradicts the first equation in 3.1.3, therefore for all implying that is not invertible.
+Note that Step 1, implies that there are at least two classes in in under relation
+- Step 2
+- Step 3
+Given a SLTM, such that or then In fact, the following list shows all the different cases for satisfying the hypothesis of this lemma, and the corresponding matrix that defines an isomorphism
+One can check in each case that is an invertible matrix and that its entries satisfy equation 2.2.6 of Theorem 2.6, (taking and ). Note that in the last case, we have since
+- Step 4
+Let be a SLTM. By Step 2 and Step 3, we have that is in the class of or in the class of Finally Step 1 implies that there are exactly two classes in
+β
+Corollary 3.4.
+If are any two matrices in If then
+Proof.
+Let be a SLTM. By the proof of Proposition 3.3, we have that:
+β’
+If then one can check that the matrix defined in Step 2 can be factored as follows:
+that is
+(3.1.4) Now this factorization yields the following well-defined process of transformations:
+(3.1.5) Is not difficult to check that and therefore one conclude that
+β’
+If then one can check that in all the cases described in Step 3, one have In fact all the corresponding matrices can be factored in a product of matrices of the form and the factorization yields a well-defined process of transformations, via operations, therefore implies that We left details to the readers.
+Let such that As a consequence of the proof of Proposition 3.3, this implies that one and only one of the following is true or Finally the above analysis implies in both cases that whenever β
 3.2 The class of
-In this section we describe a very explicit criterium (theorem 2.9) for determining when a given SLTM, is in the same class of the matrix This criterium is a generalization of lemma 2.6.
-We first need the following definition:
-Definition 2.8.
-Given a SLTM, a scalar and indices we define the triangular subdeterminant
-Note that if we take for we recover the number defined in subsection 2.2.
-Let and the generators of and respectively.
-Theorem 2.9.
-Given a SLTM, then if and only if one of the following conditions holds
-1.
-for each pair of indices
-2.
-Given a SLTM, then for each pair of indices such that we have that for each
+We would like to understand, how the information coming from the codifying matrices, can be used to describe completely an specific isomorphism class in In this direction, in this section we describe a very explicit criterion (Theorem 3.12) for determining when a given SLTM, say is in the same class of the matrix This criterion is a generalization of Step 3 in the proof of Proposition 3.3.
+Along this subsection we denote and the generators of and respectively. Also recall from subsection 1.2, the notation (or simply ) given by
 Lemma 3.5.
 If is a SLTM, such that then one of the following conditions holds
-1.
+- 1
 for each pair of indices
-2.
+or
+- 2
 For each pair of indices such that we have that for each
@@
 If is an isomorphism, then by Theorem 2.6 we have
-for each and each Lets assume that there is a for some pair with Since the matrix is invertible, there is a such that therefore equation 2.3.1 implies that
+for each and each Let us assume that there is a for some pair with Since the matrix is invertible, there is a such that therefore equation 3.2.1 implies that
 On the other hand if we take in equation 3.2.1, we can also conclude that therefore
@@
 Therefore we have
+β
 Definition 3.6.
@@
 Let a connected subgraph of We denote by the set given by:
-If the set is not empty, we say that it is a chain of leaders of If is a (not trivial) connected component of then we say that is a maximal chain of leaders.
+If the set is not empty, we say that it is a chain of leaders of If is a (non-trivial) connected component of then we say that is a maximal chain of leaders.
 If are two different connected components of then we say that and are two independent chains of leaders of
@@
 is a minimal element if and only if there is no such that and that means that the th row of is zero.
-The proof of lemma 2.12 is direct, we left to the readers.
+The proof of Lemma 3.7 is direct and it is left to the readers.
 Example 3.8.
-1.
-Let
-(2.3.6) then In this case is itself a ramified chain of leaders of The associated graph looks like this:
-Note that is composed by two simple chains: and
-2.
-Let
-(2.3.7) Then In this case the associated graph is disconnected and it looks like this:
-The sets are two independent simple chains of leaders such that
-3.
-Let
-(2.3.8) In this case then is itself a simple chain of leaders. The associated graph in this case looks like this
-4.
-Consider the matrix
-(2.3.9) In this case is itself a ramified chain, composed by two simple chains where
-The associated graph in this case looks like this:
-5.
 Consider the matrix:
-(2.3.10) Then the set is a ramified chain of leaders of The directed graph associated is given by:
-Note that if we write and then are two (not independent) simple chain of leaders of such that
-6.
-Consider the matrix:
-(2.3.11) Then the set is composed by two independent chains of leaders. The directed graph associated is given by:
-Note that is divided in two independent (maximal) chains:
-Also note that is a simple chain, while is a ramified chain, with where
+Then the set of leaders is The directed graph associated is given by:
+Note that is divided in two independent maximal chains of leaders, where
+Also note that is a simple chain, while is a ramified chain, such that where are the simple chains of leaders given by:
+Finally note that the minimal vertices of are , this is because the st, the th and the th rows of are zero row. In particular the vertex is an isolated vertex.
 Lemma 3.9.
@@
 Proof.
-Let and suppose that In particular this implies that there is a such that By our hypothesis, for each and this implies that if and only if In particular, we have that This implies that in In fact, if we have that Otherwise, if then since there is a such that and again this implies that By a recursive argument, we can find a finite sequence: such that for each that is in other words
+Let and such that By our hypothesis, for each and this implies that if and only if In particular, we have that since
+Fix an We have two cases:
 1.
-Fix an If we assume that in then we have that:
-The last equations imply that:
-Therefore we have:
+If we assume that (that is ), then by our hypothesis, we have that:
+Then we have:
+therefore
 2.
-Fix an . If there is a finite sequence: such that for each that is, then we have:
-(2.3.12) (2.3.13) (2.3.14) From the equations in 2.3.12 we obtain respectively:
-(2.3.15) Using recursively equations 2.3.13 and 2.3.14 we obtain respectively:
-(2.3.18) and
-(2.3.19) Finally we obtain:
-as desired.
+If then since then necessarily there is a such that and again this implies that By a recursive argument, we can find a finite sequence: such that for each that is (in other words ).
+In this case, by our hypothesis, we have:
+(3.2.7) (3.2.8) (3.2.9) From the equations in 3.2.7 we obtain respectively:
+(3.2.10) Using recursively equations 3.2.8 and 3.2.9 we obtain respectively:
+β
 Lemma 3.10.
-Assuming that is itself a simple chain of leaders, say such that, for each we have for each Then
+Assuming that is itself a simple chain of leaders. If for each and each we have then
 Proof.
-First note that the hypothesis implies that the presentation of is given by the relations
-also the hypothesis implies that the all the entries of are determined by the values of with Since is a chain we have that for each We also denote
-We define for each minimal element (in particular for ), the element Being minimal means that there is no such that in particular by lemma 2.12, we know that
+Recall that and denote the generators of and respectively. Let then the hypothesis implies that the presentation of is given by the relations
+Since is itself a simple chain, we have that for each We also denote We define for each minimal element (in particular for ), the element Being minimal means that there is no such that in particular by Lemma 3.7, we know that
 For we define the element recursively by:
@@
 For we assume that then
-One can check recursively (similarly to the proof of lemma 2.14) that:
-that is
-and so on.
+Recall that we want to prove that then we need to examine the two terms on the right of equation 3.2.17 to see them clearer:
+On one hand, one can check recursively (similarly to the proof of Lemma 3.9), that:
 Therefore
-and
-Now note that condition for each implies that
+On the other hand, note that condition for each implies that
 Moreover, we can use equation recursively to check that
-for any
-In fact, it is clear that
-Assuming that equation 2.3.25 is true, for any then we have:
-therefore
-and equation 2.3.25 is true for any for any
-Example 2.16.
-Let be the matrix of equation 2.3.8, then then the presentation of is given by the equations
+for any In fact, assuming that equation 3.2.20 is true, for any and noticing that that we obtain:
+Therefore equation 3.2.20 is true for any for any
 Proposition 3.11.
@@
 For each minimal element we define Again, being minimal means that there is no a such that therefore Lemma 3.7 implies that
-For not minimal, we define recursively:
-By property 2.3.4 of lemma 2.12, is well defined by equation 2.3.26.
 By item 4 of Lemma 3.7, we know that the set is a simple chain of leaders, therefore the fact that follows analogously to the proof of this in Lemma 3.10.
 Note that each independent chain can be treated independently.
-Therefore we have that the assignment defines an homomorphism
-It is clear that the associated matrix is invertible, therefore is an isomorphism. ∎
-Example 2.18.
-Example 2.19.
+Therefore we have that the assignment defines an homomorphism It is clear that the associated matrix is invertible, therefore is an isomorphism.
+β
+Theorem 3.12.
+Given a SLTM, then if and only if one of the following conditions holds
+- 1
+for each pair of indices
+or
+- 2
+For each with we have that for each
+Proof.
 Example 3.13.
 Let be the matrix of equation 3.2.6.
-As we mentioned in example 2.13 is a ramified chain, composed by two (not independent) simple chains, and One can check that satisfies all the conditions of proposition 2.17.
-The isomorphism described in the proof of proposition 2.17, is given by:
-The associated matrix is
+Note that satisfies the hypothesis of Proposition 3.11.
+As we have mentioned in Example 3.8, is divided in two independent chains, a ramified one with and a simple one
+In this case the minimal vertices of are where only is isolated. Following the proof of Proposition 3.11, we can define the isomorphism by taking:
+For minimal vertices:
+For vertices associated to chain
+For vertices associated to chain
+And finally, for vertices associated to chain
+The following Proposition provides another evidence of the veracity of Conjecture 2.21.
+Proposition 3.14.
+Let be a SLTM. If then
+Proof.
+Assuming that but We shall prove that by induction on the number of non-zero rows of
+β’
+If let be the leader of the only non-zero row of Since Theorem 3.12 implies that either or for each Both cases implies that is the only nonzero entry of and that the ETO: is well-defined over Is not difficult to see that and therefore
+β’
+If let the minimal value such that the th row of is different from zero.
+Let be the leader of the th row.
+Since Theorem 3.12 implies that either or for each Both cases implies that is the only nonzero entry of and that the ETO: is well-defined over Let then most of the entries of are equal to the corresponding entry of with the exception of those in positions of the form with such that For those cases we have the relation:
+(3.2.22) (see equation 2.3.11). Therefore in the aforementioned positions, we have that
+If is such that and is the leader of the th row of then since In particular, we have that and by definition of we have that By equation 3.2.22, we conclude that and then For any other triple with the entries of are the same as the corresponding entries of then it is clear that for each and each Theorem 3.12, implies that
+By construction and by equation 3.2.22, is a matrix with non-zero rows. By induction we conclude that and by transitivity of relation we conclude as desired.
+β
 Example 3.15.
-Let the matrix of equation 2.3.10.
-As we have mentioned in example 2.13, is a ramified chain composed by the two (maximal) simple chains: and Note that in this case the minimal vertices are all of them but are isolated vertices of Also note that satisfies the hypothesis of proposition 2.17.
-Therefore
-Following the proof of proposition 2.17, we have that:
-Using the recursive definition provided in equation 2.3.26 in the proof of proposition 2.17, we have for :
-Analogously for we obtain:
+Let be the matrix defined as the following product of elementary matrices:
+where
+One can check that the matrix defines an isomorphism and the factorization given in equation 3.2.23 yields a well-defined process of transformation via operations, that transform into This process can be deduced from the proof of Proposition 3.14.
+Note that for the isomorphism defined in example 3.13, the associated matrix can also be factored as an adequate product of elementary matrices, that yields a well-defined process of transformation via operations, that transform into We left details to the readers.
 3.3 A lower bound for the number of classes
-In this subsection we show the following:
+We would like to know, how many classes can be found in , for each and how this number depends (or not) on the ground field Let us denote the number of classes in under As we observed in Subsection 3.1, the number for is finite and do not depend on the ground field.
+Although, at this point we do not know if this number is always finite or if it is independent or not on the ground field, in this subsection we obtain a lower bound for that is still independent on the ground field:
 Theorem 3.16.
 There are at least classes in .
-To do so, we shall show that is a set of representatives of pairwise different classes under relation
-Lemma 2.24.
-Considering the matrix we have that if and only if
 Proof.
-Is not difficult to see that the matrix satisfies the conditions of theorem 2.9 if and only if ∎
-Proposition 2.25.
-The matrices and are related under if and only if
-Proof.
-By lemma 2.24, we know that whenever Therefore we only have to prove that the matrices and for are not related by
-If we assume that there is an isomorphism then by theorem 2.2, we have
+We divide the proof in several steps, where basically we shall show that the matrices are representatives of pairwise different classes.
+- Step 1
+First note that if and only if In fact, note that the matrix satisfies the first condition of Theorem 3.12, then On the other hand, if then by definition of matrix (see Subsection 1.2), the entry in position is not zero and therefore the matrix do not satisfy the necessary condition of Theorem 3.12, and then as desired.
+- Step 2
+Let and with If we assume that there is an isomorphism then by Theorem 2.6, we have
+(3.3.1) - Step 2.1
 In particular if we fix a then for each pair equation 3.3.1 reduces to
-If we assume that there is a such that then equation 2.4.2 implies that whenever and On the other hand, if we assume that for all then necessarily since But if we replace in equation 2.4.1, we obtain and therefore for all but this is a contradiction with the definition of the matrix The last analysis implies that for each there is a such that and for each with
-Now we replace in equation 2.4.1, and such that then we obtain Since by definition of the matrix then we conclude that That is valid for each Therefore, since we conclude that By the last analysis we also conclude that for each pair we have In particular, we conclude that there are exactly rows such that
-If we replace in equation 2.4.1 we obtain
-and by definition of the matrices and the indices we obtain
-Note that by the symmetry of equation 2.4.3 we can omit the hypothesis that
-Going back again to equation 2.4.1 but this time taking we obtain
-By definition of the matrix and by the facts that for each and we obtain more precisely that
-If we take for some we obtain
-Let us fix such that (we know there is at least two rows satisfying that condition. Then we have:
+(3.3.2) By our assumption, necessarily for each there is a such that (Since the associated matrix has to be invertible).
+If we assume that then equation 3.3.2 implies that whenever and On the other hand, if we assume that and we replace in equation 3.3.1, we obtain and therefore for all but this is a contradiction with the definition of the matrix The last analysis implies that for each there is a such that and for each with
+Now we replace in equation 3.3.1, and such that then we obtain Since by definition of the matrix then we conclude that That is valid for each Therefore, since we conclude that
+By the last analysis we also conclude that for each pair we have In particular, we conclude that there are exactly indices such that
+- Step 2.2
+Let us fix now a pair such that If we replace in equation 3.3.1, we obtain
+(3.3.3) and by definition of the matrices and the indices we have that whenever and whenever , therefore equation 3.3.3 reduces to
+(3.3.4) Note that by the symmetry of equation 3.3.4, we can omit the hypothesis that
+- Step 2.3
+If we replace in equation 3.3.1, we obtain
+(3.3.5) By definition of the matrix (in particular using that if and if ), and by the facts that for each and we obtain, after simplification by that
+(3.3.6) - Step 2.4
+Let us fix If we replace in equation 3.3.6, then we have that and then we obtain
+(3.3.7) - Step 2.5
+Let us fix with such that (we know there is at least two of them satisfying that condition), in particular Therefore we have the following two cases:
 - 1.
 - 2.
-The last two cases implies that implies that and then there are at most rows such that but this is a contradiction, since we know there are exactly rows satisfying that condition.
-By the last analysis we conclude that and are not isomorphic, and then ∎
-2.5 Elementary triangular operations
-In this section we define a series of matrix operations called Elementary triangular operations (ETO), that preserves the class under relation At the end of the subsection we prove for any pair that for having is necessary and sufficient that there exists of a (finite) sequence of ETO that transform into
-Definition 2.26.
-Let be a SLTM, fix an and an scalar We define the matrix as the SLTM, where
-Example 2.27.
-Consider the matrix
-then
-Lemma 2.28.
-Let ba a SLTM, fix and an scalar Then
-Proof.
-Let Let and the generators of and respectively. We claim that the assignment
-defines an isomorphism To see that recall that by equation 2.1.3, we only have to check that and the associated matrix is invertible. Now note that
-1.
-If then clearly
-2.
-If then we have
-and
-3.
-if then we have
-and
-By the last analysis, we conclude that the assignment defines an homomorphism Is not difficult to see that the associated matrix is invertible, in fact Therefore we conclude that defines an isomorphism ∎
-Definition 2.29.
-Let be a SLTM, and fix If the matrix satisfies the following conditions:
-1.
-The entry for each
-2.
-The entry for each
-then we define the matrix as the SLTM, where
-Example 2.30.
-Consider the matrix
-and take and Is not difficult to see that satisfies the conditions of definition 2.29. Then the matrix is well defined. Moreover we have
-Lemma 2.31.
-Let be a SLTM, let If satisfies the conditions of definition 2.29, then
-Proof.
-Let Let and the generators of and respectively. We claim that the assignment
-defines an isomorphism To see that, note that by our hypothesis we have:
-1.
-For we have
-and
-2.
-For we have
-and
-3.
-For it is immediate that
-4.
-For we have
-and
-5.
-For we have
-and
-The last analysis implies that the assignment defines a homomorphism Is not difficult to see that the associated matrix is invertible, in fact Therefore defines an isomorphism ∎
-Recall that for a given SLTM, we denote
-Definition 2.32.
-Let be a SLTM, take an scalar and assume that there are indices such that for each If one of the following conditions holds:
-1.
-2.
-and for each
-Then we define the matrix as the SLTM, given by
-Example 2.33.
-Consider the matrix
-Let and Note that the conditions of definition 2.32 hold. In this case Therefore we have
-Example 2.34.
-Therefore we have
-Lemma 2.35.
-Let a SLTM, let be a pair of indices and an scalar. Assuming that the conditions of definition 2.32 hold, then we have that
-Proof.
-Let Let and the generators of and respectively. We divide the analysis in two cases:
-1.
-If then we have that the matrix is given by:
-We claim that the assignment:
-defines an isomorphism In fact, if then it is clear that
-We also can check for that
-and
-Finally for we have
-and
-Therefore, the assignment defines a morphism Is not difficult to see that the matrix is invertible, moreover then is an isomorphism.
-2.
-If and for each then the matrix as the SLTM, is given by
-We claim that the assignment:
-defines an isomorphism In fact, if then it is clear that
-We also can check for that
-On the other hand
-Note that by our hypothesis, we have then
-then as desired.
-Finally for we have
-and
-Therefore, the assignment defines a morphism Is not difficult to see that the matrix is invertible, moreover then is an isomorphism.
-Definition 2.36.
-We call Elementary triangular operations (ETO for short) any of the following matrix assignments (whenever they are well defined):
-Given two SLTM, and we say that they are equivalent by ETO if there is a finite sequence of ETO that transform into We denote by whenever and are equivalent by ETO.
-Example 2.37.
-Consider the matrices
-then In fact one can check that
-Definition 2.38.
-In the following we shall denote respectively by the matrices given by:
-Note that the matrix described in the proof of lemma 2.28 is of the form the matrix described in the proof of lemma 2.31 is of the form and the matrix described in the proof of lemma 2.35 is of the form
-Theorem 2.39.
-Let be two SLTM, then if and only if
-Proof.
-By lemmas 2.28, 2.31 and 2.35 we have that if then Moreover lemmas 2.28, 2.31 and 2.35 implies that for two SLTM, we have that if and only if there is an invertible matrix that can be written as a (finite) product of matrices of the form and
-If then there exist an invertible matrix satisfying equation 2.1.6.
-Since is invertible, it can be factored as a product of elementary matrices (in the usual sense), Now the result follows from the fact that the matrices and are elementary matrices and any other elementary matrix can be written as a product of matrices of the form and ∎
-3 Applications To Soergel Calculus
-Fix a Coxeter system of finite rank (with for certain set of indices ), and a field where the number has sense in for each pair Following [6], the geometric realization of over a is a triple composed by a vector space subsets and where:
-and the equations:
-define representations of the group on and respectively. For any and any let
-We are avoiding a technical definition on the diagrammatic Soergel category since it is well known.
-We invite the readers to look for more information in the original work of Elias and Williamson [3].
-In this section we shall assume that the ground field has characteristic different from and satisfies all the technicalities described in [3].
-Let the free monoid generated by the symbols In [12], for each element we define the SLTM:
-and the corresponding nil graded algebra We have proved in [12], that the algebra corresponds to the Gelfand-Tsetlin subalgebra of the endomorphism algebra in the diagrammatic Soergel Category of Elias and Williamson (see [3]), relative to the Libedynski’s light leaves basis (see [8]) and the set of Jucys-Murphy elements defined by Ryom-Hansen in [19].
-In [12] we show that for the case of a Coxeter system of type the algebras are invariant under commuting moves, and we proved that by defining explicit isomorphisms.
-Now in theorem 3.2 we generalize this fact for any other type.
-In this case we use the techniques developed in this article, to obtain a cleaner proof.
-Let us recall first, certain technical definitions coming from [12]:
-Definition 3.1.
-1.
-Let we define the column vector associated to as the matrix given by:
-2.
-Given two elements of say and we define the matrix by
-where
-In [12], we show that for any composed expression we have that:
-Moreover in [12], we proved that for we have the following (kind of) associative property:
-where
-Now we are ready to prove the following:
-Theorem 3.2.
-For any Coxeter system the algebra is invariant under commuting moves.
-Proof.
-Let such that that its, the elements and of the group commute. Let and then are elements of that differ by a commuting move.
-1.
-Let and Then and Therefore:
-(3.0.3) where
-2.
-Note that by definition and Therefore we have:
-By the same reason, we have that two first rows of are given by:
-analogously the two first rows of are given by:
-3.
-Note that for every we have that Let us denote
-then
-4.
-Following with the notation of the s, we have that the rows of the matrices from the third to the last one are given respectively by:
-and
-But since and commute, we clearly have that
-- 5.
-Theorem 3.2 naturally opens the following question:
-- Q3:
-Is the algebra invariant under other braid moves?
-Fortunately we have an answer for Q3, although it is not a positive one:
-Proposition 3.3.
-The algebra is not invariant under braid moves.
-Proof.
-It is enough to provide an example where two equivalent by braid moves expressions and generate two not isomorphic algebras and
-Let be the Coxeter system of type Let and then:
-1.
-It is clear that and are equivalent by braid moves.
-2.
-The corresponding matrices and are given by:
-3.
-Considering the matrices defined by:
-one can check that they are well defined (see subsection for details) and by definition More specifically we have:
-4.
-Let be the entries of let be the entries of the matrix By theorem 2.39, if the algebras and were isomorphic, then necessarily there will be an isomorphism Assuming that, then the associated matrix would satisfy equation 2.1.6 of theorem 2.2.
-In particular we would have for and any pair that:
-(3.0.4) -
-5.
-Equation 3.0.4 in particular implies that:
-- 6.
-7.
-Let and be the sets of generators of the algebras and respectively.
-From the last two items, we conclude that if there exists an isomorphism then necessarily for In particular this implies that the matrices and are in the same class under relation but this contradicts theorem 2.9.
-In fact
-The analysis above implies that and are not isomorphic. ∎
+The last two cases imply that whenever we necessarily have Then there are at most indices such that but this is a contradiction, since we know from Step 2.1, that there are exactly of them satisfying that condition (recall that ).
+By the last analysis we conclude that and are not isomorphic, and then whenever
+- Step 3
+By Step 1 and Step 2, we conclude that the matrices are representatives of pairwise different classes under relation Therefore there are at least different classes in .
+β
+- Q1:
+Is the number always finite?
+- Q2:
+Is the number independent of the ground field?
+We do not have an answer for this questions yet, but motivated by our intuition, we define the following:
+Conjecture 3.17.
+The number for depends on the ground field. Moreover the number is finite if and only if the ground field is finite.
+We expect to be able to prove or disprove Conjecture 3.17 in future works.